OpenStudy (anonymous):

A helium–neon laser emits a beam of circular cross section with a radius r and a power P.? (a) Find the maximum electric field in the beam. (Use any variable or symbol stated above along with the following as necessary: μ0 and c. (b) What total energy is contained in a length of the beam? (Use the following as necessary: P,l , and c.)(c) Find the momentum carried by a length of the beam. (Use the following as necessary: P, l, and c.)

OpenStudy (anonymous):

there is a relation between intensity of light beam and the magnitude of electric field.\[I = (1/2) c \epsilon _{o} n E^{2} = P/\pi r^{2}\] \[E ^{2} = 2P/ c \epsilon_{o} n \pi r^{2}\] E= magnitude of electric field n= refractive index of medium \[\mu_{o}\epsilon_{0} = 1/c^{2}\] put the values and get the answer

OpenStudy (anonymous):

energy= power*time = P*(1m/speed of light)\[energy= (P*1m)/c \]

OpenStudy (anonymous):

momentum= energy/c